# Strategic Content Creation with GenAI: To Share or Not to Share?

### v2

@@ C.4 Strong Monotonicity and Lipschitz Continuity of ESE
We’ll extend the monotonicity result from Lemma˜3 to show that with $\beta$-strongly convex cost functions, the game becomes strongly monotone.
@@
  • From Lemma 15 and the arguments in the proof of Theorem 5, we know that $\mathcal{X}_{i}(\rho)>0$ for all $i\in[n]$. Consequently, terms like $\left\|\mathbf{x}\right\|^{\gamma-2}$ are well-defined and $C^{1}$ in a neighborhood of $\mathcal{X}(\rho)$ and are continuously differentiable as a composition of $C^{1}$ functions.
Next, we need to verify that the Jacobian matrix $J_{x}(\mathbf{x},\rho)=\frac{\partial\mathbf{g}}{\partial\mathbf{x}}$ is invertible at $\mathcal{X}(\rho)$. Proposition 27 establishes that when the cost functions $c_{i}(x_{i})$ are $\beta$-strongly convex, the enforced sharing game is strongly monotone. Specifically, for any distinct $\mathbf{x},\mathbf{x}^{\prime}$ in the strategy space, the gradient map $\mathbf{g}(\mathbf{x},\rho)$ satisfies:
@@
C.5 Proofs of Section 4’s Main Results
@@
The proof is structured in two main parts: First, establishing the approximation guarantees of Algorithm 1, and second, analyzing its computational complexity under an efficient implementation.
@@
We now analyze the computational complexity of an efficient implementation of Algorithm 1. The algorithm performs $\lceil 1/\delta\rceil$ iterations, where in each iteration it computes an approximate ESE and checks if it constitutes an $\eta$-FSE.
For computing the approximate ESE (Line 6), we leverage Proposition 27, which states that under the assumption $\forall i\in[n],x_{i}\in[0,x_{i}^{\max}]:c^{\prime\prime}_{i}(x_{i})\geq\beta$, the enforced sharing game is strongly monotone. For strongly monotone games, the Multi-Agent Mirror Descent (MAMD) algorithm with full gradient feedback converges at a rate of $O(1/T)$, as established in [12] (see Algorithm˜2 for a full specification of the algorithm). Therefore, to find a $\delta$-approximation of an ESE (w.r.t. the $l_{2}$ norm), it suffices to run MAMD for $T=O(1/\delta)$ iterations.
@@
For the $\eta$-FSE check (Line 7), we need to verify if any player can profitably deviate from the profile $(\mathbf{x},\mathbf{1})$ by more than $\eta$. By Proposition 2, for any fixed quality choice $x_{i}$, the optimal sharing level is either $s_{i}=0$ or $s_{i}=1$. Therefore, for each player $i$, we need to check two types of deviations: 1) Deviations where $s_{i}=1$ (changing only quality); and 2) Deviations where $s_{i}=0$ (changing quality and not sharing).
@@
5:  Compute $L_{0}$ using Lemma 26
@@
Step 2: Analyzing the constants in Theorem 8
We now analyze the constants that appear in the proof of Theorem 8. Starting with the constant $A$, we have:
@@
Appendix D Omitted Proofs from Section 6
@@ D.3 Proofs for Alternative Optimization Objectives
Replacing Lemma 25 with this bound, the proof of Theorem˜8 extends directly. Lemma˜24 remains unchanged, while the objective discretization error becomes at most $(L_{f}+L_{\rho})(1+L_{\text{ESE}})\delta$, where $L_{\text{ESE}}=B(0,1/\beta,0.5,1,2-\gamma)$ is the Lipschitz constant of $\mathcal{X}(\cdot)$ from Lemma˜28. Consequently, choosing $\delta=\varepsilon/A^{\prime}$ with
@@ D.4 Proofs for GenAI-Dependent Traffic
$\tau_{i}=\frac{1}{\alpha}\left[\left(\frac{\left\|\mathbf{x}\right\|+\alpha\sum_{j=1}^{n}x_{j}s_{j}}{\left\|\mathbf{x}\right\|+\alpha\sum_{j\neq i}x_{j}s_{j}}\right)^{1-\gamma}-1\right].$

### v4

@@ C.4 Strong Monotonicity and Lipschitz Continuity of ESE
We extend the monotonicity result from Lemma˜3 to show that with $\beta$-strongly convex cost functions, the game becomes strongly monotone.
@@
  • From Lemma˜15 and the arguments in the proof of Theorem˜5, we know that $\mathcal{X}_{i}(\rho)>0$ for all $i\in[n]$. Consequently, terms like $\left\|\mathbf{x}\right\|^{\gamma-2}$ are well-defined and $C^{1}$ in a neighborhood of $\mathcal{X}(\rho)$ and are continuously differentiable as a composition of $C^{1}$ functions.
Next, we need to verify that the Jacobian matrix $J_{x}(\mathbf{x},\rho)=\frac{\partial\mathbf{g}}{\partial\mathbf{x}}$ is invertible at $\mathcal{X}(\rho)$. Lemma˜27 establishes that when the cost functions $c_{i}(x_{i})$ are $\beta$-strongly convex, the enforced sharing game is strongly monotone. Specifically, for any distinct $\mathbf{x},\mathbf{x}^{\prime}$ in the strategy space, the gradient map $\mathbf{g}(\mathbf{x},\rho)$ satisfies:
@@
C.5 Proofs of Section˜4’s Main Results
@@
The proof is structured in two main parts: First, establishing the approximation guarantees of Algorithm˜1, and second, analyzing its computational complexity under an efficient implementation.
@@
We now analyze the computational complexity of an efficient implementation of Algorithm˜1. The algorithm performs $\lceil 1/\delta\rceil$ iterations, where in each iteration it computes an approximate ESE and checks if it constitutes an $\eta$-FSE.
For computing the approximate ESE (Line 6), we leverage Lemma˜27, which states that under the assumption $\forall i\in[n],x_{i}\in[0,x_{i}^{\max}]:c^{\prime\prime}_{i}(x_{i})\geq\beta$, the enforced sharing game is strongly monotone. For strongly monotone games, the Multi-Agent Mirror Descent (MAMD) algorithm with full gradient feedback converges at a rate of $O(1/T)$, as established in [12] (see Algorithm˜2 for a full specification of the algorithm). Therefore, to find a $\delta$-approximation of an ESE (w.r.t. the $l_{2}$ norm), it suffices to run MAMD for $T=O(1/\delta)$ iterations.
@@
For the $\eta$-FSE check (Line 7), we need to verify if any player can profitably deviate from the profile $(\mathbf{x},\mathbf{1})$ by more than $\eta$. By Lemma˜2, for any fixed quality choice $x_{i}$, the optimal sharing level is either $s_{i}=0$ or $s_{i}=1$. Therefore, for each player $i$, we need to check two types of deviations: 1) Deviations where $s_{i}=1$ (changing only quality); and 2) Deviations where $s_{i}=0$ (changing quality and not sharing).
@@
5:  Compute $L_{0}$ using Lemma˜26
@@
Step 2: Analyzing the constants in Theorem˜8
We now analyze the constants that appear in the proof of Theorem˜8. Starting with the constant $A$, we have:
@@
Appendix D Omitted Proofs from Section˜6
@@ D.3 Proofs for Alternative Optimization Objectives
Replacing Lemma˜25 with this bound, the proof of Theorem˜8 extends directly. Lemma˜24 remains unchanged, while the objective discretization error becomes at most $(L_{f}+L_{\rho})(1+L_{\text{ESE}})\delta$, where $L_{\text{ESE}}=B(0,1/\beta,0.5,1,2-\gamma)$ is the Lipschitz constant of $\mathcal{X}(\cdot)$ from Lemma˜28. Consequently, choosing $\delta=\varepsilon/A^{\prime}$ with
@@ D.4 Proofs for GenAI-Dependent Traffic
$\tau_{i}=\frac{1}{\alpha}\left[\left(\frac{\left\|\mathbf{x}\right\|+\alpha\left(x_{i}+\sum_{j\neq i}^{n}x_{j}s_{j}\right)}{\left\|\mathbf{x}\right\|+\alpha\sum_{j\neq i}x_{j}s_{j}}\right)^{1-\gamma}-1\right].$